OpenStudy (anonymous):

:Give an example of two matrices of order 2x2 , with all the elements non-zeros, such that AB=0 where A,B are the two matrices

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 1 \\ 1 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

and\[\left[\begin{matrix}-1 & -1 \\ 1 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

yw